OpenStudy (anonymous):

5 [lim x ^3/2. = 5 (4)^3/2. = 40. How? I can't get the numbers to add up.

OpenStudy (agent0smith):

\[\LARGE 5 \times \lim_{x \rightarrow 4} x ^{\frac{ 3 }{ 2 }} = \] Is this what you have...?

OpenStudy (agent0smith):

\[\LARGE 5 \times 4^{\frac{ 3 }{ 2 }} = 5 \times (4^{\frac{ 1 }{ 2 }})^3 = \] \[ \LARGE = 5 \times (\sqrt4)^3\]

OpenStudy (anonymous):

Ah! I see that now :)

OpenStudy (anonymous):

I was doing cubed root 4 to power 2

OpenStudy (agent0smith):

you can also do it as \[\LARGE 5 \times 4^{\frac{ 3 }{ 2 }} = 5 \times \sqrt{4^3} = 5 \times \sqrt{64}\]

OpenStudy (anonymous):

ok, so how does that = 64??

OpenStudy (anonymous):

duh never mind. my blonbe roots are showing :)

OpenStudy (agent0smith):

haha no prob :) it's easier if you do it the first way (sqrt of 4 first, then cubed).